Circular No. 5974

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

COMET TAKAMIZAWA-LEVY (1994f) Kesao Takamizawa and David H. Levy report their discoveries of what

appears to be the same comet, as follows:

1994 UT  α_{2000}  δ_{2000}  m_{1} Observer
Apr. 14.756  21^{h}24.^{m}0  +4^{o}48′  12.5 Takamizawa
15.4  21 23.0  +5 45  10.5 Levy

K. Takamizawa (Saku-cho, Minami-Saku-gun, Nagano-ken, Japan). Two 4-min photographic films, 0.10-m f/4.0 camera. Comet is diffuse with central condensation and coma diameter ∼ 1′. Communicated by S. Nakano.

D. H. Levy (Tucson, AZ). Visual observation with 0.41-m reflector. Comet diffuse with slight condensation and a broad, faint, fan-shaped tail. Motion toward northwest at ∼ 0^{o}.3/day.

SUPERNOVA 1994J IN ANONYMOUS GALAXY B. Schmidt and R. Kirshner, Harvard-Smithsonian Center for Astro-

physics (CfA), report that a spectrum (320–900 nm) taken by J. Huchra (CfA) with the Multiple Mirror Telescope (+ Blue Channel) on Apr. 12 UT shows this object to be of type Ia, ∼ 3 months past maximum light. The redshift of the host galaxy, derived from Ca H+K absorption and Hα emission, is cz ≈ 0.06, in fair agreement with that reported by Filippenko et al. (IAUC 5971).

GRO J0422+32 C. Chevalier and S. A. Ilovaisky, Observatoire de Haute-Provence,

write: “CCD photometry of the optical counterpart of this x-ray transient has been obtained with the 1.2-m OHP telescope in the quiescent state during two runs: Feb. 1–9 and Mar. 3–10, mostly in the I_{c} and R_{c} bands. Variance analysis of this new set shows power at a period of 5.09 hr (consistent with the value previously derived from our 1993 data; see IAUC 5692) and also at one-half, one-third, and even one-fourth of this value. Folding the data reveals a complex and variable light curve which is far from being of an ellipsoidal type. The maximum full amplitude is 0.2 mag, but it varies from night to night. The mean V magnitude for both runs was 20.66 and 20.56, respectively, with V − I_{c} = +1.83 in February.”

1994 April 15 Daniel W. E. Green